A Complete Bibliography of Publications in Acta Numerica

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

03 July 2025 Version 1.21

Title word cross-reference

3 [Hou09]. l_{1} [NN13]. qd [Par95]. R [Sha07].

-functions [Sha07].

/nuclear [NN13].

accelerated [AAD^{+}16]. Acceleration [Saa25]. accuracy [BL07]. Accurate [DDHK08]. acoustic [CWGLS12]. Acta [Hig96]. Adaptive [BCNV24, EEHJ95, Ode18]. adaptively [LGB11]. Adaptivity [BHR09]. Adjoint [GS02]. algebra [AAD^{+}16, BCD^{+}14, Chu08, DHvdV93, DDHK08, Eld06, HM22, KV17, MT20]. Algebraic [XZ17, Mär92, MU23, WS11].

Algorithms [Noc92, Pow98, BCD^{+}14, CM94, Chu08, HM22, KV17, MT20, MS06, NN13, OST20, Par95, PFM19, QMV17, Set96, Woh11]. analyses [Cuc16]. Analysis [Ell94, GS02, AM94, Bru04, Cha01, CST00, HS19, Le 05, NSZ17, DM24, Slo92, Sma97, Stu94, Wan10, dVBSV14, vDKS93]. analytic [Hou09, Sha07]. ANU [Ano92b, Ano92a, Ano93b, Ano93a, Ano94b, Ano94a, Ano95b, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano09b, Ano09a, Ano10b, Ano10a, Ano11b, Ano11a, Ano12b, Ano12a, Ano13b, Ano13a, Ano14b, Ano14a, Ano15b, Ano15a, Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a, Ano22b, Ano22a, Ano23b, Ano23a, Ano24b, Ano24a]. application [CST00]. Applications [Las24, AFW06, Flo15, Gau96, MU23, QMV17, Set96]. approach [BBO03, BR01, Hou09]. approximate [GL94, GL95]. Approximation [BIV95, CD15, Pin99, PFM19, Bof10, DeV98, DHP21, Tem08, Wat98]. approximations [Tad03]. arithmetic [BJMM23, MS06, Rum10]. art [Coo97]. Aspects [Ell94, AM94, EE01, Tem07]. assimilation [Rei19]. Asymptotic [ES08, Jin22, CWGLS12]. Asymptotic-preserving [Jin22]. atmospheric [Cul07]. Atomistic [LO13]. Atomistic-to-continuum [LO13]. Automatic [Hen96, Gri03]. averages [BL07].

Back [Ano92a, Ano93a, Ano94a, Ano95a, Ano96a, Ano97a, Ano98a, Ano99a, Ano09a, Ano10a, Ano11a, Ano12a, Ano13a, Ano14a, Ano15a, Ano16a, Ano17a, Ano18a, Ano19a, Ano20a, Ano21a, Ano22a, Ano23a, Ano24a]. barrier [SKH23]. barycentric [Flo15]. based [Fre03, GW21]. bases [Ban96]. basis [Buh00, FF15, HPR22]. Bayesian [Stu10]. behind [Coo97]. beyond [AHP21, GZ22b]. bifurcation [CST00]. Binary [FZ10]. Blow [Hou09]. Blow-up [Hou09]. Boltzmann [NS95]. Boundary [Hou95, CWGLS12, Hag99, Pes02, Slo92]. bounds [BCD^{+}14].

calculation [Bjö04]. Calculations [Pow98]. calculus [AFW06, Hac14]. Cambridge [Hig96]. cancer [DSW12]. cardiovascular [QMV17]. Carlo [BRSS18, Caf98, DKS13, Gil15]. challenges [Cha01, CT99]. chemistry [Le 05]. Choice [Wat98]. circulation [Hig06]. clinical [QMV17]. cloud [Car14]. combinatorial [PFM19]. Communication [BCD^{+}14]. Compatible [Cot23]. Complete [Neu04]. complex [GGMR09]. Complexity [Sma97]. composite [GM94]. Computation [DDE05, Tre99, FMT16, Gau96, Lus96, Mis02, Tad07, WZZ21]. Computational [ER03, Le 05, Rin97, AM94, CT99, Ede00, Hip02, Hou09, JMS11, SNCH04]. computations [HJM24, Lim21]. computed [BL07]. Computing [HAM10, LL20, AAD^{+}16, DvdS12, PFM19]. condition [Cuc16]. conditions [Hag99]. cones [TV23]. conjugate [MS06]. conservation [MS07, Tad03]. consistent [Woh11]. Constrained [Wri92]. constraint [Neu04]. Constructing [Coo97]. contact [HS19, Woh11]. Continuation [AG93, Li97]. Continuous [Coc03, CP16, Neu04]. continuum [LO13]. Control [MU23, BR01, GZ22b]. controllability [GL94, GL95]. controlled [LM14]. convection [Sty05]. convection-diffusion [Sty05]. convergence [Yse93]. coordinates [Flo15]. coupling [LO13]. Cover [Ano92b, Ano92a, Ano93b, Ano93a, Ano94b, Ano94a, Ano95b, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano09b, Ano09a, Ano10b, Ano10a, Ano11b, Ano11a, Ano12b, Ano12a, Ano13b, Ano13a, Ano14b, Ano14a, Ano15b, Ano15a, Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a, Ano22b, Ano22a, Ano23b, Ano23a, Ano24b, Ano24a]. crystalline [Lus96]. crystals [WZZ21]. cubature [Coo97]. cuisine [LCBM12]. current [Cha01]. curvature [DDE05]. Cut [BHLZ25].

D [Hou09]. Data [Heg01, Rei19, AMÖS19, Car14, Eld06, GW21, GWZ14, Wat98]. data-driven [AMÖS19]. decomposition [CM94]. Deep [BMR21, Bel21, DSW12]. denoising [LCBM12]. density [LLY19]. dependence [Coc03]. dependent [HPR22, KL98, LM14, Tad03]. Derivative [LMW19]. Derivative-free [LMW19]. design [HJM24]. deterministic [Pir12]. devices [Rin97]. difference [Tad03]. different [Kre92]. Differential [EEHJ95, Bac23, BMZG09, BCM24, Bru04, DDE05, FS94, GWZ14, KL98, LS16, Mär92, MU23, MN96, PJY97, Pla99]. differential-algebraic [MU23]. differentiation [Gri03]. diffusion [Sty05]. dimensional [CD15, DKS13, GGMR09, SG11]. dimensions [GR97]. Direct [Pow98, DRSL16, GGMR09]. discontinuous [DG25]. Discrete [MW01]. discretization [CMKO11, Woh11]. discretizations [SG11]. dissipation [LM14]. distributed [GL94, GL95]. Distributionally [KSW25]. Domain [CM94, GZ22a]. domains [GGMR09]. driven [AMÖS19]. Duality [GS02]. dynamical [Chu08, Stu94]. dynamics [BL07, Cot23, LL20, LS16, Mis02, SKH23].

efficient [DDHK08, Mis02]. Eigenvalue [LO96, Bof10, CG02, GT17, Sor02]. eigenvalues [Ips98]. elastic [Cia01]. electromagnetism [Hip02]. electrostatic [GM94]. element [AFW06, BBO03, Ban96, BR01, BBMR23, Bof10, BCNV24, BHLZ25, Cha01, Cot23, DE13, For93, GWZ14]. elements [Hip02]. Ensemble [CRS25]. Entropy [Tad03]. Equation [NS95, GR97]. Equations [EEHJ95, Bac23, BMZG09, BCM24, Bru04, Dah97, DDE05, DP14, FS94, For93, GGMR09, GWZ14, Gut97, Hou09, JMS11, Kel18, KL98, Kur18, LS16, Mär92, MN96, PJY97, Pla99]. Errata [Ano95c]. Error [GS02, Slo92, BR01, Coc03]. Essentially [Shu20]. estimation [BR01, Can06, Coc03]. Euler [Hou09]. evaluation [DDHK08, GM94]. Exact [GL94, GL95]. experimental [HJM24]. Exponential [HO10]. expression [DDHK08]. exterior [AFW06].

Fast [GGMR09, OST20, GR97]. fear [Bel21]. field [CRS25]. fields [GM94]. Filters [Tad07]. Finite [AFW06, Bof10, DE13, For93, Hip02, Kur18, MS07, BBO03, Ban96, BR01, BCNV24, BHLZ25, Cha01, Cot23, GWZ14, LGB11, MS06]. Finite-volume [Kur18]. first [NN13]. first-order [NN13]. Fit [Bel21]. fitting [Wat98]. fixed [Saa25]. fixed-point [Saa25]. Floating [BJMM23, Rum10]. Floating-point [BJMM23, Rum10]. flow [DDE05]. flows [Cul07]. fluid [CST00, Cot23]. following [AG93]. Formalization [AM94]. formulae [Coo97]. Formulations [HJM24]. Foundations [MT20]. fractional [DDG^{+}20]. Free [Hou95, LMW19]. frequency [CWGLS12, ER03]. Front [Ano92b, Ano93b, Ano94b, Ano95b, Ano96b, Ano97b, Ano98b, Ano99b, Ano09b, Ano10b, Ano11b, Ano12b, Ano13b, Ano14b, Ano15b, Ano16b, Ano17b, Ano18b, Ano19b, Ano20b, Ano21b, Ano22b, Ano23b, Ano24b]. function [Wat98]. functional [BMZG09, Bru04, LLY19]. functions [Buh00, FF15, HAM10, Sha07, Tem07]. futures [Pir12].

Galerkin [DG25]. General [But06].

Generalized [Flo15, BBO03]. generation [Hen96]. Geometric [BRSS18, EE01, HLW03, DDE05, DL02]. geometry [Com24, Ede00, Sab94, Sha07, WS11]. geophysical [Cot23]. Gibbs [Tad07]. global [Neu04]. governed [HK25]. gradient [MS06]. Greedy [Tem08]. grid [Hen96]. grids [BHR09, BG04]. group [IMKNZ00].

Hamiltonian [BRSS18, MU23, SS92]. heart [SNCH04]. hemivariational [HS19]. heterogeneous [AWEVE12]. Hierarchical [Ban96]. hierarchy [Las24]. High [DKS13, DvdS12, CWGLS12, CD15, ER03, SG11]. High-dimensional [DKS13, CD15, SG11]. high-frequency [CWGLS12]. High-performance [DvdS12]. highly [PJY97]. homogeneous [TV23]. homogenization [AHP21, ES08]. homological [AFW06]. homotopy [Li97]. hyperbolic [GWZ25, MS07]. hyperthermia [DSW12].

illustrated [HLW03]. image [AM94, LCBM12, TO05]. imaging [CP16]. immersed [Pes02]. incompressible [Hou09]. inequalities [Can06, HS19]. informed [DM24]. initial [vDKS93]. input [GWZ14]. integer [BKL^{+}13]. integral [Bru04, CWGLS12, GGMR09, Slo92]. integration [DKS13, HLW03]. integrators [BRSS18, HO10, MW01, SS92]. integro [Bru04]. integro-differential [Bru04]. interfaces [Set96]. Interior [NT08, Wri92]. Interior-point [NT08]. interpolation [Bel21]. Introduction [EEHJ95, CP16, Pla99]. Inverse [Stu10, AMÖS19, BB18, CG02, GW21]. isogeometric [dVBSV14]. issue [Ano92b, Ano92a, Ano93b, Ano93a, Ano94b, Ano94a, Ano95b, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano09b, Ano09a]. iterations [Saa25]. Iterative [FGN92, MN96].

Kalman [CRS25]. Kepler [Wan10]. Kernel [SW06]. kinematics [WS11]. kinetic [DP14]. Kohn [LLY19]. Krylov [EE01, Fre03].

Lanczos [Gut97, MS06]. Lanczos-type [Gut97]. Laplace [GR97]. large [AAD^{+}16, GOT05, Sor02]. large-scale [AAD^{+}16, GOT05]. laws [MS07, Tad03]. Learning [GW21, BMR21, Bel21, DM24, SW06, SKH23]. least [Bjö04, ST25]. least-squares [ST25]. level [Set96, TO05].

Lie [IMKNZ00]. Lie-group [IMKNZ00]. Linear [AAD^{+}16, Chu08, FGN92, TV23, vDKS93, BCD^{+}14, Bjö04, But06, DRSL16, DHvdV93, DDHK08, Eld06, Gut97, HM22, KV17, MT20, MN96, ST25]. linearly [Cia01]. liquid [WZZ21]. Low [Bac23]. Low-rank [Bac23]. lower [BCD^{+}14].

machine [DM24, SW06, SKH23]. machines [FZ10]. materials [GM94]. Mathematical [Cia01, DSW12, JMS11, QMV17, dVBSV14, Bel21, Gri03]. matrix [ER05, HAM10, Ips98, TV23]. matter [Ano92b, Ano92a, Ano93b, Ano93a, Ano94b, Ano94a, Ano95b, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano09b, Ano09a, Ano10b, Ano10a, Ano11b, Ano11a, Ano12b, Ano12a, Ano13b, Ano13a, Ano14b, Ano14a, Ano15b, Ano15a, Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a, Ano22b, Ano22a, Ano23b, Ano23a, Ano24b, Ano24a]. mean [CRS25, DDE05]. mean-field [CRS25]. measure [FMT16]. measure-valued [FMT16]. mechanics [CST00, HS19, MW01]. meshes [Ede00]. meshless [BBO03, SW06]. method [AWEVE12, Ban96, BBMR23, BRSS18, DG25, HLW03, Pes02, GR97]. Methods [EEHJ95, GS02, NS95, Wri92, BBO03, Bac23, BR01, BB18, BCM24, BCNV24, BHLZ25, But06, Caf98, CRS25, CWGLS12, Coc03, Com24, Cot23, DDG^{+}20, Dah97, DRSL16, DP14, DE13, EE01, FS94, Fre03, GZ22a, Gil15, GOT05, GWZ14, HPR22, IMKNZ00, JMS11, Kel18, LMW19, LM14, LGB11, LS16, Li97, LLY19, Mär92, MQ02, MS07, Nat99, NT08, Pir12, Pla99, Rin97, Rum10, Saa25, SW06, Set96, Slo92, Sor02, TO05, XZ17, Yse93, dVBSV14]. microstructure [Lus96]. minimization [NN13]. mining [Eld06, Heg01]. Mixed [BKL^{+}13, HM22]. Mixed-integer [BKL^{+}13]. MLP [Pin99]. Model [Fre03, GW21, Pin99]. Modelling [Cul07, WZZ21, Ben21, Cia01, DL02, Hig06, LGB11, Ode18, QMV17, SNCH04]. models [AMÖS19, DDG^{+}20, GW21, DM24]. Modern [BB18, Can06]. Molecular [BL07, LS16, SKH23]. mollifiers [Tad07]. Moment [Las24]. Moment-SOS [Las24]. monotone [Com24]. Monte [BRSS18, Caf98, DKS13, Gil15]. motion [ÖVBS17]. moving [BHR09]. multicore [AAD^{+}16]. multigrid [XZ17, Yse93]. Multilevel [Gil15]. Multipole [GR97]. Multiscale [SNCH04, AWEVE12, Dah97, Jin22, Ode18]. Multivariate [dB93, Li97].

Navier [For93, Hou09]. network [DHP21]. Networks [Ell94, Pin99, DM24]. Neural [DHP21, Ell94, Pin99, DM24]. Newton [Wan10]. no [Hou09]. non [Shu20]. non-oscillatory [Shu20]. Nonlinear [DeV98, BKL^{+}13, GOT05, GT17, Kel18, NSZ17, Tad03]. nonlocal [DDG^{+}20]. nonsymmetric [Gut97]. norm [NN13]. norms [Wat98]. nuclear [NN13]. numbers [Cuc16]. Numerica [Hig96]. Numerical [AHP21, BGL05, CWGLS12, CT99, DDG^{+}20, DP14, Eld06, Ell94, GOT05, Hac14, HS19, Hig06, Hou95, Kel18, LM14, Li97, LLY19, Mär92, Nat99, NSZ17, PJY97, DM24, Sab94, Sor02, Stu94, Tem07, WS11, BCD^{+}14, BMZG09, Ben21, Bru04, CST00, DHvdV93, ES08, GM94, Hag99, HLW03, HM22, KV17, Le 05, MT20, Pla99, QMV17, Sma97, Wan10, Woh11, vDKS93]. Numerical-asymptotic [CWGLS12]. numerically [BL07]. numerics [FS20].

ocean [Hig06]. Old [Yse93]. operator [Com24, Dah97]. operators [SKH23]. Optimal [Ben21, HJM24, BCD^{+}14, BR01, GZ22b]. Optimization [HK25, LO96, Noc92, Pow98, Tod01, Wri92, BKL^{+}13, CP16, GOT05, KSW25, LMW19, NT08, Neu04, TV23]. oracle [Can06]. order [NN13]. ordinary [PJY97]. Orthogonal [Gau96, OST20]. oscillatory [PJY97, Shu20]. outlook [DvdS12]. Overcoming [SKH23]. overview [SS92].

Padé [BIV95]. parabolic [GWZ25]. Parallel [DHvdV93]. parallelization [GWZ25]. parameter [GL94, GL95]. parametric [CD15, SG11]. Partial [LS16, Bac23, DDE05, FS94, GWZ14]. Particle [NS95]. path [AG93]. pattern [Car14]. PDEs [CD15, DE13, FF15, GZ22b, GS02, HK25, NSZ17, SG11]. performance [DvdS12]. perspective [CRS25, Le 05, Rei19, Stu10]. perspectives [GW21]. perturbation [Ips98]. Petrov [DG25]. phase [FS20]. phenomena [Bel21]. phenomenon [Tad07]. physical [Jin22]. physics [GW21, DM24]. physics-based [GW21]. physics-informed [DM24]. piecewise [dB93]. planning [DSW12]. point [BGL05, BJMM23, Car14, NT08, Rum10, Saa25]. polynomial [Li97]. polynomials [Gau96, OST20, dB93]. port [MU23]. port-Hamiltonian [MU23]. posteriori [BR01, GS02]. Postprocessing [GS02]. precision [HM22, MS06]. Preconditioning [Wat15]. predictive [Ode18]. Preface [Ano92c]. preserving [CMKO11, Jin22]. Press [Hig96]. Pricing [Pir12]. principles [SKH23]. prism [Bel21]. Probabilistic [Cuc16]. problem [GT17]. Problems [Hou95, Kre92, AMÖS19, BB18, BGL05, Bjö04, Bof10, CG02, CST00, GWZ25, GW21, HK25, HPR22, Jin22, SS92, ST25, Sor02, Stu10, Sty05, Tad03, Woh11, vDKS93]. Programming [BT95]. proofs [Yse93]. propagating [Set96]. propagation [ER03]. pseudospectra [Tre99]. pseudospectral [FS94].

Quadratic [BT95]. quantum [LL20, Rin97]. quasi [Caf98, DKS13]. quasi-Monte [Caf98, DKS13].

Radial [Buh00, FF15]. Radiation [Hag99]. Random [ER05, GWZ14]. Randomized [KV17, MT20]. rank [Bac23]. recognition [Car14]. Reduced [HPR22]. reduction [Fre03, GW21]. refined [LGB11]. regime [LL20]. regional [DSW12]. regularization [BB18]. related [Las24, DM24, Tad03]. Relative [Ips98]. relativity [CT99]. remarkable [Bel21]. ResNets [GZ22b]. results [Cha01, Ips98, Rum10]. retarded [BMZG09]. retrieval [FS20]. Review [Hig96, FS94]. Rigorous [Rum10, Mis02]. robust [KSW25].

saddle [BGL05]. satisfaction [Neu04]. scale [AAD^{+}16, AHP21, GOT05, LM14]. scales [Kre92]. scattering [CWGLS12]. schemes [DL02, Jin22, Kur18, Shu20, Woh11]. Schrödinger [JMS11, Rei19]. Schwarz [GZ22a]. science [CT99, Coo97, TO05]. scientific [PFM19]. Search [Pow98, Neu04]. Secrets [LCBM12]. Section [Ano95c]. Semi [Sha07]. Semi-analytic [Sha07]. semiclassical [JMS11, LL20, Rin97]. semiconductor [Rin97]. Semidefinite [Tod01]. separation [AHP21, FZ10]. Sequential [BT95]. set [Set96, TO05]. shallow [Kur18]. shallow-water [Kur18]. Sham [LLY19]. shells [Cha01, Cia01]. shocks [LM14]. simulation [Ben21, Hag99]. singular [Ips98]. small [LM14]. small-scale [LM14]. software [AAD^{+}16]. Solution [FGN92, BMZG09, BGL05, For93, Li97, MN96, PJY97, vDKS93]. Solutions [Hou95, FMT16]. solvers [GGMR09, Gut97]. Solving [AMÖS19, FF15, FS94]. Some [Cha01]. SOS [Las24]. Sparse [BG04, SG11, ST25, DRSL16]. special [Tem07]. Splitting [BCM24, MQ02, Com24]. squares [Bjö04, ST25]. Stability [KL98, Tad03, vDKS93]. state [Sty05]. statistical [BMR21, Can06]. Status [DvdS12]. Steady [Sty05]. Steady-state [Sty05]. Stochastic [GWZ14, LS16, Pla99, SG11]. Stokes [For93, Hou09]. Störmer [HLW03]. strongly [NSZ17]. structure [CMKO11, ÖVBS17]. structure-preserving [CMKO11]. Structured [CG02]. Subdivision [DL02]. subspace [EE01]. subspaces [Fre03]. support [FZ10]. surface [DE13]. surfaces [Sab94]. Survey [BBO03, DRSL16, ÖVBS17]. Symplectic [SS92]. system [QMV17]. Systems [FGN92, Chu08, DRSL16, DvdS12, GL94, GL95, Gut97, HK25, Li97, MU23, MN96, Stu94].

Taste [BIV95]. techniques

[AFW06, Heg01, Mis02, SW06]. tensor [Bac23, Hac14, SG11]. Tensors [Lim21]. Theory [Noc92, Set96, ER05, EE01, LLY19, Pin99, Sma97, Tad03]. therapy [DSW12]. three [GR97, GGMR09]. three-dimensional [GGMR09]. Time [GWZ25, HPR22, Kre92, KL98, Tad03]. time-dependent [HPR22, KL98, Tad03]. timescale [SKH23]. tomography [Nat99]. Topics [CMKO11, Las24]. Topological [Car14, Mis02]. Total [TO05]. training [FZ10]. Transfer [SKH23]. transport [Rin97]. transportation [Ben21]. trends [BMZG09]. Triangulations [Ede00]. truncation [GZ22a]. Tsunami [LGB11]. Turnpike [GZ22b]. type [Bru04, Gut97].

uncertainties [HK25]. Unconstrained [Noc92]. unified [BBO03, Hou09]. University [Hig96]. using [AMÖS19, OST20, Rum10].

value [vDKS93]. valued [FMT16]. values [Ips98]. variation [TO05]. variational [MW01, SKH23, dVBSV14]. Variationally [Woh11]. vector [FZ10]. Verification [Rum10]. Verlet [HLW03]. version [GR97]. via [Can06]. view [Gri03]. viewpoint [BMR21]. virtual [BBMR23]. viscosity [Coc03]. Volterra [Bru04]. volume [Ano92b, Ano92a, Ano93b, Ano93a, Ano94b, Ano94a, Ano95b, Ano95a, Ano96b, Ano96a, Ano97b, Ano97a, Ano98b, Ano98a, Ano99b, Ano99a, Ano09b, Ano09a, Ano10b, Ano10a, Ano11b, Ano11a, Ano12b, Ano12a, Ano13b, Ano13a, Ano14b, Ano14a, Ano15b, Ano15a, Ano16b, Ano16a, Ano17b, Ano17a, Ano18b, Ano18a, Ano19b, Ano19a, Ano20b, Ano20a, Ano21b, Ano21a, Ano22b, Ano22a, Ano23b, Ano23a, Ano24b, Ano24a, Kur18, LGB11, MS07].

water [Kur18]. wave [ER03]. Wavelet [Dah97]. Wavelets [DL92]. waves [Hag99]. way [DKS13]. weighted [Shu20]. without [Bel21].

References

Abdelfattah:2016:LAS

[AAD^{+}16] A. Abdelfattah, H. Anzt, J. Dongarra, M. Gates, A. Haidar, J. Kurzak, P. Luszczek, S. Tomov, I. Yamazaki, and A. YarKhan. Linear algebra software for largescale accelerated multicore computing. Acta Numerica, 25:1– 160, 2016. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Arnold:2006:FEE

[AFW06] Douglas N. Arnold, Richard S.

Falk, and Ragnar Winther. Finite element exterior calculus, homological techniques, and applications. Acta Numerica, 15:1– 155, 2006. CODEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Allgower:1993:CPF

[AG93] Eugene L. Allgower and Kurt Georg. Continuation and path following. Acta Numerica, 2:1– 64, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Altmann:2021:NHB

[AHP21] Robert Altmann, Patrick Henning, and Daniel Peterseim. Numerical homogenization beyond scale separation. Acta Numerica, 30:1–86, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
numerical-homogenization-
beyond-scale-separation/03C5F8C9AFB7E2180C05A9E8DBD8F525.

Alvarez:1994:FCA

[AM94] Luis Alvarez and Jean Michel Morel. Formalization and computational aspects of image analysis. Acta Numerica, 3:1–59, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Arridge:2019:SIP

[AMÖS19] Simon Arridge, Peter Maass, Ozan Öktem, and Carola-Bibiane Schönlieb. Solving inverse problems using datadriven models. Acta Numerica, 28:1–174, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
solving-inverse-problems-
using-datadriven-models/CE5B3725869AEAF46E04874115B0AB15.

Anonymous:1992:AVIb

[Ano92a] Anonymous. ANU volume 1 issue 1 cover and back matter. Acta Numerica, 1:b1, 1992. CO-DEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1992:AVIa

[Ano92b] Anonymous. ANU volume 1 issue 1 cover and front matter. Acta Numerica, 1:f1–f6, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1992:P

[Ano92c] Anonymous. Preface. Acta Numerica, 1:i, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1993:AVIb

[Ano93a] Anonymous. ANU volume 2 issue 1 cover and back matter. Acta Numerica, 2:b1–b2, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1993:AVIa

[Ano93b] Anonymous. ANU volume 2 issue 1 cover and front matter. Acta Numerica, 2:f1–f6, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1994:AVIb

[Ano94a] Anonymous. ANU volume 3 issue 1 cover and back matter. Acta Numerica, 3:b1, 1994. CO-DEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1994:AVIa

[Ano94b] Anonymous. ANU volume 3 issue 1 cover and front matter. Acta Numerica, 3:f1–f6, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1995:AVIb

[Ano95a] Anonymous. ANU volume 4 issue 1 cover and back matter. Acta Numerica, 4:b1, 1995. CO-DEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1995:AVIa

[Ano95b] Anonymous. ANU volume 4 issue 1 cover and front matter. Acta Numerica, 4:f1–f6, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1995:ES

[Ano95c] Anonymous. Errata for section 1.10.7. Acta Numerica, 4:329– 333, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1996:AVIb

[Ano96a] Anonymous. ANU volume 5 issue 1 cover and back matter. Acta Numerica, 5:b1, 1996. CO-DEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1996:AVIa

[Ano96b] Anonymous. ANU volume 5 issue 1 cover and front matter. Acta Numerica, 5:f1–f6, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1997:AVIb

[Ano97a] Anonymous. ANU volume 6 issue 1 cover and back matter. Acta Numerica, 6:b1, 1997. CO-DEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1997:AVIa

[Ano97b] Anonymous. ANU volume 6 issue 1 cover and front matter. Acta Numerica, 6:f1–f6, 1997. CODEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1998:AVIb

[Ano98a] Anonymous. ANU volume 7 issue 1 cover and back matter. Acta Numerica, 7:b1, 1998. CO-DEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1998:AVIa

[Ano98b] Anonymous. ANU volume 7 issue 1 cover and front matter. Acta Numerica, 7:f1–f6, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1999:AVIb

[Ano99a] Anonymous. ANU volume 8 issue 1 cover and back matter. Acta Numerica, 8:b1, 1999. CO-DEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:1999:AVIa

[Ano99b] Anonymous. ANU volume 8 issue 1 cover and front matter. Acta Numerica, 8:f1–f6, 1999. CODEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2009:AVIb

[Ano09a] Anonymous. ANU volume 18 issue 1 cover and back matter. Acta Numerica, 18:b1, 2009. CODEN ANUMFU. ISBN 0-521-19211-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2009:AVIa

[Ano09b] Anonymous. ANU volume 18 issue 1 cover and front matter. Acta Numerica, 18:f1–f6, 2009. CODEN ANUMFU. ISBN 0-521-19211-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2010:AVCb

[Ano10a] Anonymous. ANU volume 19 cover and back matter. Acta Numerica, 19:b1, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2010:AVCa

[Ano10b] Anonymous. ANU volume 19 cover and front matter. Acta Numerica, 19:f1–f6, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2011:AVCb

[Ano11a] Anonymous. ANU volume 20 cover and back matter. Acta Numerica, 20:b1, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2011:AVCa

[Ano11b] Anonymous. ANU volume 20 cover and front matter. Acta Numerica, 20:f1–f6, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2012:AVCb

[Ano12a] Anonymous. ANU volume 21 cover and back matter. Acta Numerica, 21:b1, 2012. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2012:AVCa

[Ano12b] Anonymous. ANU volume 21 cover and front matter. Acta Numerica, 21:f1–f6, 2012. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2013:AVCb

[Ano13a] Anonymous. ANU volume 22 cover and back matter. Acta Numerica, 22:b1, 2013. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2013:AVCa

[Ano13b] Anonymous. ANU volume 22 cover and front matter. Acta Numerica, 22:f1–f6, 2013. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2014:AVCb

[Ano14a] Anonymous. ANU volume 23 cover and back matter. Acta Numerica, 23:b1, 2014. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2014:AVCa

[Ano14b] Anonymous. ANU volume 23 cover and front matter. Acta Numerica, 23:f1–f6, 2014. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2015:AVCb

[Ano15a] Anonymous. ANU volume 24 cover and back matter. Acta Numerica, 24:b1–b2, 2015. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2015:AVCa

[Ano15b] Anonymous. ANU volume 24 cover and front matter. Acta Numerica, 24:f1–f8, 2015. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2016:AVCb

[Ano16a] Anonymous. ANU volume 25 cover and back matter. Acta Numerica, 25:b1–b2, 2016. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2016:AVCa

[Ano16b] Anonymous. ANU volume 25 cover and front matter. Acta Numerica, 25:f1–f8, 2016. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Anonymous:2017:AVCb

[Ano17a] Anonymous. ANU volume 26 cover and back matter. Acta Numerica, 26:b1–b2, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
product/CD47E20B56033D9B928B91880B69CCCC.

Anonymous:2017:AVCa

[Ano17b] Anonymous. ANU volume 26 cover and front matter. Acta Numerica, 26:f1–f8, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
product/CAB2088CCC091A35910D54485B702CA7.

Anonymous:2018:AVCb

[Ano18a] Anonymous. ANU volume 27 cover and back matter. Acta Numerica, 27:b1–b2, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-27-cover-and-back-
matter/4D99F47E66773F0D8987D931D1CC41B7.

Anonymous:2018:AVCa

[Ano18b] Anonymous. ANU volume 27 cover and front matter. Acta Numerica, 27:f1–f7, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-27-cover-and-front-
matter/31ED918D5D0E3C645E4759F83B66E1D2.

Anonymous:2019:AVCb

[Ano19a] Anonymous. ANU volume 28 cover and back matter. Acta Numerica, 28:b1–b2, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-28-cover-and-back-
matter/37BC3C8988AFA1AF421F91EB8A24BD09.

Anonymous:2019:AVCa

[Ano19b] Anonymous. ANU volume 28 cover and front matter. Acta Numerica, 28:f1–f7, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-28-cover-and-front-
matter/50D5848124876F27AC959618E401EED4.

Anonymous:2020:AVCb

[Ano20a] Anonymous. ANU volume 29 cover and back matter. Acta Numerica, 29:b1–b2, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-29-cover-and-back-
matter/2E6DF7A8CF6A034A8696FE2BCC4F8271.

Anonymous:2020:AVCa

[Ano20b] Anonymous. ANU volume 29 cover and front matter. Acta Numerica, 29:f1–f6, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        anu-volume-29-cover-and-front-
                                        matter/B9B842CEB66A178E4EC235A8AC6EE3D3.

Anonymous:2021:AVCb

[Ano21a] Anonymous. ANU volume 30 cover and back matter. Acta Numerica, 30:b1, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        anu-volume-30-cover-and-back-
                                        matter/4FF8C6A1D40BA56A770C11BB6CEFAFFF.

Anonymous:2021:AVCa

[Ano21b] Anonymous. ANU volume 30 cover and front matter. Acta Numerica, 30:f1–f6, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        anu-volume-30-cover-and-front-
                                        matter/285F45741F1792F77EEFE0CF7450CA12.

Anonymous:2022:AVCb

[Ano22a] Anonymous. ANU volume 31 cover and back matter. Acta Numerica, 31:b1, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        anu-volume-31-cover-and-back-
                                        matter/010752884B6EC8AB88909AB2F3778E50.

Anonymous:2022:AVCa

[Ano22b] Anonymous. ANU volume 31 cover and front matter.

Acta Numerica, 31:f1–f6, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-31-cover-and-front-
matter/0E919884CCB419E7AF48770120082D5D.

Anonymous:2023:AVCb

[Ano23a] Anonymous. ANU volume 32 cover and back matter. Acta Numerica, 32:b1, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-32-cover-and-back-
matter/14164335F9751B8D3D760A5BAC8586A2.

Anonymous:2023:AVCa

[Ano23b] Anonymous. ANU volume 32 cover and front matter. Acta Numerica, 32:f1–f6, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-32-cover-and-front-
matter/6562E29E2EB7BDEF62786728707EBE98.

Anonymous:2024:AVCb

[Ano24a] Anonymous. ANU volume 33 cover and back matter. Acta Numerica, 33:b1, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
anu-volume-33-cover-and-back-
matter/7D6D690AB00BFB07D758E83C4049BD2C.

Anonymous:2024:AVCa

[Ano24b] Anonymous. ANU volume 33 cover and front matter. Acta Numerica, 33:f1–f6, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        anu-volume-33-cover-and-front-
                                        matter/64A52A5A8439A7814D1B825CE7552676.

Abdulle:2012:HMM

[AWEVE12] Assyr Abdulle, E. Weinan, Björn Engquist, and Eric Vanden-Eijnden. The heterogeneous multiscale method. Acta Numerica, 21:1–87, 2012. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Bachmayr:2023:LRT

[Bac23] Markus Bachmayr. Low-rank tensor methods for partial differential equations. Acta Numerica, 32:1–121, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        lowrank-tensor-methods-for-
                                        partial-differential-equations/
                                        06D2B93A3E4FB412CB13F8B107F399AB.

Bank:1996:HBF

[Ban96] Randolph E. Bank. Hierarchical bases and the finite element method. Acta Numerica, 5:1–43, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Benning:2018:MRM

[BB18] Martin Benning and Martin Burger. Modern regularization methods for inverse problems. Acta Numerica, 27:1– 111, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
modern-regularization-methods-
for-inverse-problems/1C84F0E91BF20EC36D8E846EF8CCB830.

BeiraoDaVeiga:2023:VEM

[BBMR23] Lourenço Beirão Da Veiga, Franco Brezzi, L. Donatella Marini, and Alessandro Russo. The virtual element method. Acta Numerica, 32:123–202, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
virtual-element-method/F02A0BB9FFFF84C511AE6F9BDB2660C1.

Babuska:2003:SMG

[BBO03] Ivo Babuška, Uday Banerjee, and John E. Osborn. Survey of meshless and generalized finite element methods: a unified approach. Acta Numerica, 12:1– 125, 2003. CODEN ANUMFU. ISBN 0-521-82523-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ballard:2014:CLB

[BCD^{+}14] G. Ballard, E. Carson, J. Demmel, M. Hoemmen, N. Knight, and O. Schwartz. Communication lower bounds and optimal algorithms for numerical linear algebra. Acta Numerica, 23:1– 155, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Blanes:2024:SMD

[BCM24] Sergio Blanes, Fernando Casas, and Ander Murua. Splitting methods for differential equations. Acta Numerica, 33:1–161, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
splitting-methods-for-differential-
equations/2C1BD434F8D2052E593A755351B3ACA3.

Bonito:2024:AFE

[BCNV24] Andrea Bonito, Claudio Canuto, Ricardo H. Nochetto, and Andreas Veeser. Adaptive finite element methods. Acta Numerica, 33:163–485, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
adaptive-finite-element-methods/
5D36C8F8DBE998FDFDF7E49179898B11.

Belkin:2021:FFR

[Bel21] Mikhail Belkin. Fit without fear: remarkable mathematical phenomena of deep learning through the prism of interpolation. Acta Numerica, 30:203–248, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/

journals/acta-numerica/article/
fit-without-fear-remarkable-
mathematical-phenomena-of-
deep-learning-through-the-
prism-of-interpolation/DBAC769EB7F4DBA5C4720932C2826014.

Benamou:2021:OTM

[Ben21] Jean-David Benamou. Optimal transportation, modelling and numerical simulation. Acta Numerica, 30:249–325, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
optimal-transportation-modelling-
and-numerical-simulation/
D6D857468618B8EEE6F85FE197564E69.

Bungartz:2004:SG

[BG04] Hans-Joachim Bungartz and Michael Griebel. Sparse grids. Acta Numerica, 13:147–269, 2004. CODEN ANUMFU. ISBN 0-521-83811-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Benzi:2005:NSS

[BGL05] Michele Benzi, Gene H. Golub, and Jörg Liesen. Numerical solution of saddle point problems. Acta Numerica, 14:1–137, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Burman:2025:CFE

[BHLZ25] Erik Burman, Peter Hansbo, Mats G. Larson, and Sara Zahedi. Cut finite element methods. Acta Numerica, 34:1–121, July 2025. CODEN ANUMFU.

ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        cut-finite-element-methods/
                                        97D22C4D003C93666739D1BCF7AEBAF7.

Budd:2009:AMG

[BHR09] Chris J. Budd, Weizhang Huang, and Robert D. Russell. Adaptivity with moving grids. Acta Numerica, 18:111–241, 2009. CO-DEN ANUMFU. ISBN 0-521-19211-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Brezinski:1995:TPA

[BIV95] C. Brezinski, Ufr Ieea, and J. Van Iseghem. A taste of Padé approximation. Acta Numerica, 4:53–103, 1995. CODEN ANUMFU. ISBN 0-521-48255-

0. ISSN 0962-4929 (print), 1474-

0508 (electronic).

Boldo:2023:FPA

[BJMM23] Sylvie Boldo, Claude-Pierre Jeannerod, Guillaume Melquiond, and Jean-Michel Muller. Floatingpoint arithmetic. Acta Numerica, 32:203–290, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        floatingpoint-arithmetic/
                                        287C4D5F6D4A43FBEEB1ABED2A405AAF.

Bjorck:2004:CLL

[Bjö04] Åke Björck. The calculation of linear least squares problems. Acta Numerica, 13:1–53, 2004.

CODEN ANUMFU. ISBN 0-521-83811-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Belotti:2013:MIN

[BKL^{+}13] Pietro Belotti, Christian Kirches, Sven Leyffer, Jeff Linderoth, James Luedtke, and Ashutosh Mahajan. Mixed-integer nonlinear optimization. Acta Numerica, 22:1–131, 2013. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Bond:2007:MDA

[BL07] Stephen D. Bond and Benedict J. Leimkuhler. Molecular dynamics and the accuracy of numerically computed averages. Acta Numerica, 16:1–65, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Bartlett:2021:DLS

[BMR21] Peter L. Bartlett, Andrea Montanari, and Alexander Rakhlin. Deep learning: a statistical viewpoint. Acta Numerica, 30:87–201, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
deep-learning-a-statistical-
viewpoint/7BCB89D860CEDDD5726088FAD64F2A5A.

Bellen:2009:RTN

[BMZG09] Alfredo Bellen, Stefano Maset, Marino Zennaro, and Nicola Guglielmi. Recent trends in the numerical solution of retarded functional differential equations. Acta Numerica, 18:1–110, 2009. CODEN ANUMFU. ISBN 0-521-51642-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Boffi:2010:FEA

[Bof10] Daniele Boffi. Finite element approximation of eigenvalue problems. Acta Numerica, 19:1–120, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Becker:2001:OCA

[BR01] Roland Becker and Rolf Rannacher. An optimal control approach to a posteriori error estimation in finite element methods. Acta Numerica, 10:1–102, 2001. CODEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Bou-Rabee:2018:GIH

[BRSS18] Nawaf Bou-Rabee and J. M. Sanz-Serna. Geometric integrators and the Hamiltonian Monte Carlo method. Acta Numerica, 27:113–206, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        geometric-integrators-and-
                                        the-hamiltonian-monte-carlo-
                                        method/3319BBAC1CF3EA2BE160ADB2D74FD1C4.

Brunner:2004:NAF

[Bru04] Hermann Brunner. The numerical analysis of functional integral and integro-differential equations of Volterra type. Acta Numerica, 13:55–145, 2004. CO-DEN ANUMFU. ISBN 0-521-83811-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Boggs:1995:SQP

[BT95] Paul T. Boggs and Jon W. Tolle.

Sequential quadratic programming. Acta Numerica, 4:1–51, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Buhmann:2000:RBF

[Buh00] M. D. Buhmann. Radial basis functions. Acta Numerica, 9:1–38, 2000. CODEN ANUMFU. ISBN 0-521-78037-

3. ISSN 0962-4929 (print), 1474-0508 (electronic). URL http:

//journals.cambridge.org/
action/displayAbstract?fromPage=
online&aid=68665.

Butcher:2006:GLM

[But06] J. C. Butcher. General linear methods. Acta Numerica, 15:157–256, 2006. CODEN ANUMFU. ISBN 0-521-86815-

7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Caflisch:1998:MCQ

[Caf98] Russel E. Caflisch. Monte Carlo and quasi-Monte Carlo methods. Acta Numerica, 7:1–49, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Candes:2006:MSE

[Can06] Emmanuel J. Candès. Modern statistical estimation via oracle inequalities. Acta Numerica, 15:257–325, 2006. CODEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Carlsson:2014:TPR

[Car14] Gunnar Carlsson. Topological pattern recognition for point cloud data. Acta Numerica, 23:289–368, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cohen:2015:AHD

[CD15] Albert Cohen and Ronald De-Vore. Approximation of highdimensional parametric PDEs. Acta Numerica, 24:1–159, 2015. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Chu:2002:SIE

[CG02] Moody T. Chu and Gene H. Golub. Structured inverse eigenvalue problems. Acta Numerica, 11:1–71, 2002. CO-DEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Chapelle:2001:SNR

[Cha01] Dominique Chapelle. Some new results and current challenges in the finite element analysis of shells. Acta Numerica, 10:215– 250, 2001. CODEN ANUMFU.

ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Chu:2008:LAA

[Chu08] Moody T. Chu. Linear algebra algorithms as dynamical systems. Acta Numerica, 17:1–86, 2008. CODEN ANUMFU. ISBN 0-521-51642-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ciarlet:2001:MML

[Cia01] Philippe G. Ciarlet. Mathematical modelling of linearly elastic shells. Acta Numerica, 10:103– 214, 2001. CODEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Chan:1994:DDA

[CM94] Tony F. Chan and Tarek P.

Mathew. Domain decomposition algorithms. Acta Numerica, 3:61–143, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Christiansen:2011:TSP

[CMKO11] Snorre H. Christiansen, Hans Z.

Munthe-Kaas, and Brynjulf Owren. Topics in structurepreserving discretization. Acta Numerica, 20:1–119, 2011. CO-DEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cockburn:2003:CDE

[Coc03] Bernardo Cockburn. Continuous dependence and error estimation for viscosity methods. Acta Numerica, 12:127–180, 2003. CO-DEN ANUMFU. ISBN 0-521-82523-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Combettes:2024:GMO

[Com24] Patrick L. Combettes. The geometry of monotone operator splitting methods. Acta Numerica, 33:487–632, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
geometry-of-monotone-operator-
splitting-methods/AC76B29B4F3B229F93810C1FCB35102C.

Cools:1997:CCF

[Coo97] Ronald Cools. Constructing cubature formulae: the science behind the art. Acta Numerica, 6: 1–54, 1997. CODEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cotter:2023:CFE

[Cot23] Colin J. Cotter. Compatible finite element methods for geophysical fluid dynamics. Acta Numerica, 32:291–393, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
compatible-finite-element-
methods-for-geophysical-fluid-
dynamics/D2A87A43005D82261B9C7C4BB7CD000E.

Chambolle:2016:ICO

[CP16] Antonin Chambolle and Thomas Pock. An introduction to continuous optimization for imaging. Acta Numerica, 25:161– 319, 2016. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Calvello:2025:EKM

[CRS25] Edoardo Calvello, Sebastian Reich, and Andrew M. Stuart. Ensemble Kalman methods: a mean-field perspective. Acta Numerica, 34:123–291, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
ensemble-kalman-methods-a-
meanfield-perspective/94C9B874BBD4F11B8D36DD42D9F01BC7.

Cliffe:2000:NAB

[CST00] K. A. Cliffe, A. Spence, and S. J. Tavener. The numerical analysis of bifurcation problems with application to fluid mechanics. Acta Numerica, 9:39–131, 2000. CODEN ANUMFU. ISBN 0-521-78037-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cook:1999:NRC

[CT99] Gregory B. Cook and Saul A.

Teukolsky. Numerical relativity: challenges for computational science. Acta Numerica, 8:1–45, 1999. CODEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cucker:2016:PAC

[Cuc16]  Felipe Cucker.  Probabilis-

tic analyses of condition numbers. Acta Numerica, 25:321– 382, 2016. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Cullen:2007:MAF

[Cul07] Mike Cullen. Modelling atmospheric flows. Acta Numerica, 16:67–154, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Chandler-Wilde:2012:NAB

[CWGLS12] Simon N. Chandler-Wilde, Ivan G. Graham, Stephen Langdon, and Euan A. Spence. Numericalasymptotic boundary integral methods in high-frequency acoustic scattering. Acta Numerica, 21:89–305, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dahmen:1997:WMM

[Dah97] Wolfgang Dahmen. Wavelet and multiscale methods for operator equations. Acta Numerica, 6:55– 228, 1997. CODEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

deBoor:1993:MPP

[dB93] C. de Boor. Multivariate piecewise polynomials. Acta Numerica, 2:65–109, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Deckelnick:2005:CGP

[DDE05] Klaus Deckelnick, Gerhard Dziuk, and Charles M. Elliott. Computation of geometric partial differential equations and mean curvature flow. Acta Numerica, 14:139–232, 2005. CO-DEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

DElia:2020:NMN

[DDG^{+}20] Marta D’Elia, Qiang Du, Christian Glusa, Max Gunzburger, Xiaochuan Tian, and Zhi Zhou. Numerical methods for nonlocal and fractional models. Acta Numerica, 29:1–124, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
numerical-methods-for-nonlocal-
and-fractional-models/C5B2FDDA1C929337B8DAFD04CD219D85.

Demmel:2008:AEE

[DDHK08] James Demmel, Ioana Dumitriu, Olga Holtz, and Plamen Koev. Accurate and efficient expression evaluation and linear algebra. Acta Numerica, 17:87–145, 2008. CODEN ANUMFU. ISBN 0-521-51642-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dziuk:2013:FEM

[DE13] Gerhard Dziuk and Charles M.

Elliott. Finite element methods for surface PDEs. Acta Numerica, 22:289–396, 2013. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

DeVore:1998:NA

[DeV98] Ronald A. DeVore. Nonlinear approximation. Acta Numerica, 7:51–150, 1998. CODEN ANUMFU. ISBN 0-521-64316-

3. ISSN 0962-4929 (print), 1474-

0508 (electronic).

Demkowicz:2025:DPG

[DG25] Leszek Demkowicz and Jay Gopalakrishnan. The discontinuous Petrov–Galerkin method. Acta Numerica, 34:293–384, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        discontinuous-petrovgalerkin-
                                        method/71BCF32CDE92B0924051FA31E8F54DC2.

DeVore:2021:NNA

[DHP21] Ronald DeVore, Boris Hanin, and Guergana Petrova. Neural network approximation. Acta Numerica, 30:327–444, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        neural-network-approximation/
                                        7077A90FB36D405D903DCC82683B7A48.

Demmel:1993:PNL

[DHvdV93] James W. Demmel, Michael T. Heath, and Henk A. van der Vorst. Parallel numerical linear algebra. Acta Numerica, 2:111– 197, 1993. CODEN ANUMFU.

ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dick:2013:HDI

[DKS13] Josef Dick, Frances Y. Kuo, and Ian H. Sloan. High-dimensional integration: The quasi-Monte Carlo way. Acta Numerica, 22:133–288, 2013. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

DeVore:1992:W

[DL92] Ronald A. DeVore and Bradley J.

Lucier. Wavelets. Acta Numerica, 1:1–56, 1992. CO-DEN ANUMFU. ISBN 0-521-41026-6. ISSN 0962-4929 (print), 1474-0508 (electronic). URL ftp://ftp.

gwdg.de/pub/math/wavelets/
papers/waveletGeneral.ps.
gz.

Dyn:2002:SSG

[DL02] Nira Dyn and David Levin.

Subdivision schemes in geometric modelling. Acta Numerica, 11:73–144, 2002. CO-DEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Ryck:2024:NAP

[DM24] Tim De Ryck and Siddhartha Mishra. Numerical analysis of physics-informed neural networks and related models in physics-informed machine learning. Acta Numerica, 33:633–713, July 2024. CODEN ANUMFU.

ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        numerical-analysis-of-physicsinformed-
                                        neural-networks-and-related-
                                        models-in-physicsinformed-
                                        machine-learning/A059C6E13478F0F7C70EC7C976716F9F.

Dimarco:2014:NMK

[DP14] G. Dimarco and L. Pareschi. Numerical methods for kinetic equations. Acta Numerica, 23:369–520, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Davis:2016:SDM

[DRSL16] Timothy A. Davis, Sivasankaran Rajamanickam, and Wissam M. Sid-Lakhdar. A survey of direct methods for sparse linear systems. Acta Numerica, 25:383– 566, 2016. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Deuflhard:2012:MCT

[DSW12] Peter Deuflhard, Anton Schiela, and Martin Weiser. Mathematical cancer therapy planning in deep regional hyperthermia. Acta Numerica, 21:307– 378, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

daVeiga:2014:MAV

[dVBSV14] L. Beirão da Veiga, A. Buffa, G. Sangalli, and R. Vázquez. Mathematical analysis of variational isogeometric methods. Acta Numerica, 23:157–287, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dongarra:2012:HPC

[DvdS12] J. J. Dongarra and A. J. van der Steen. High-performance computing systems: Status and outlook. Acta Numerica, 21:379– 474, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Edelsbrunner:2000:TMC

[Ede00] Herbert Edelsbrunner. Triangulations and meshes in computational geometry. Acta Numerica, 9:133–213, 2000. CODEN ANUMFU. ISBN 0-521-78037-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Eiermann:2001:GAT

[EE01] Michael Eiermann and Oliver G.

Ernst. Geometric aspects of the theory of Krylov subspace methods. Acta Numerica, 10:251–312, 2001. CODEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Eriksson:1995:IAM

[EEHJ95] Kenneth Eriksson, Don Estep, Peter Hansbo, and Claes Johnson. Introduction to adaptive methods for differential equations. Acta Numerica, 4:105– 158, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Elden:2006:NLA

[Eld06] Lars Eldén. Numerical linear algebra in data mining. Acta Numerica, 15:327–384, 2006. CO-DEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ellacott:1994:ANA

[Ell94] S. W. Ellacott. Aspects of the numerical analysis of neural networks. Acta Numerica, 3:145– 202, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Engquist:2003:CHF

[ER03] Björn Engquist and Olof Runborg. Computational high frequency wave propagation. Acta Numerica, 12:181–266, 2003. CODEN ANUMFU. ISBN 0-521-82523-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Edelman:2005:RMT

[ER05] Alan Edelman and N. Raj Rao. Random matrix theory. Acta Numerica, 14:233–297, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Engquist:2008:ANH

[ES08] B. Engquist and P. E. Souganidis. Asymptotic and numerical homogenization. Acta Numerica, 17:147–190, 2008. CODEN ANUMFU. ISBN 0-521-51642-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Fornberg:2015:SPR

[FF15] Bengt Fornberg and Natasha Flyer. Solving PDEs with radial basis functions. Acta Numerica, 24:215–258, 2015. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Freund:1992:ISL

[FGN92] Roland W. Freund, Gene H.

Golub, and Noël M. Nachtigal. Iterative solution of linear systems. Acta Numerica, 1:57–100, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic). Cited in [?].

Floater:2015:GBC

[Flo15] Michael S. Floater. Generalized barycentric coordinates and applications. Acta Numerica, 24:161–214, 2015. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Fjordholm:2016:CMV

[FMT16] Ulrik S. Fjordholm, Siddhartha Mishra, and Eitan Tadmor. On the computation of measurevalued solutions. Acta Numerica, 25:567–679, 2016. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Fortin:1993:FES

[For93] Michel Fortin. Finite element solution of the Navier–Stokes equations. Acta Numerica, 2:239– 284, 1993. CODEN ANUMFU.

ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Freund:2003:MRM

[Fre03] Roland W. Freund. Model reduction methods based on Krylov subspaces. Acta Numerica, 12:267–319, 2003. CODEN ANUMFU. ISBN 0-521-82523-

7. ISSN 0962-4929 (print), 1474-0508 (electronic). URL http:

//cm.bell-labs.com/cm/cs/
doc/03/4-01.ps.gz.

Fornberg:1994:RPM

[FS94] Bengt Fornberg and David M. Sloan. A review of pseudospectral methods for solving partial differential equations. Acta Numerica, 3:203–267, 1994. CO-DEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Fannjiang:2020:NPR

[FS20] Albert Fannjiang and Thomas Strohmer. The numerics of phase retrieval. Acta Numerica, 29:125–228, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
numerics-of-phase-retrieval/
7000D4BBF45E088E39268D88BE49F730.

Fletcher:2010:BST

[FZ10] Roger Fletcher and Gaetano Zanghirati. Binary separation and training support vector machines. Acta Numerica, 19:121– 158, 2010. CODEN ANUMFU.

ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Gautschi:1996:OPA

[Gau96] Walter Gautschi. Orthogonal polynomials: applications and computation. Acta Numerica, 5:45–119, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Greengard:2009:FDS

[GGMR09] Leslie Greengard, Denis Gueyffier, Per-Gunnar Martinsson, and Vladimir Rokhlin. Fast direct solvers for integral equations in complex three-dimensional domains. Acta Numerica, 18:243– 275, 2009. CODEN ANUMFU. ISBN 0-521-19211-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Giles:2015:MMC

[Gil15] Michael B. Giles. Multilevel Monte Carlo methods. Acta Numerica, 24:259–328, 2015. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Glowinski:1994:EAC

[GL94] R. Glowinski and J. L. Lions. Exact and approximate controllability for distributed parameter systems. Acta Numerica, 3:269– 378, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Glowinski:1995:EAC

[GL95] R. Glowinski and J. L. Lions. Exact and approximate controllability for distributed parameter systems. Acta Numerica, 4:159– 328, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Greengard:1994:NEE

[GM94] Leslie Greengard and Monique Moura. On the numerical evaluation of electrostatic fields in composite materials. Acta Numerica, 3:379–410, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Gould:2005:NML

[GOT05] Nick Gould, Dominique Orban, and Philippe Toint. Numerical methods for large-scale nonlinear optimization. Acta Numerica, 14:299–361, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Greengard:1997:NVF

[GR97] Leslie Greengard and Vladimir Rokhlin. A new version of the Fast Multipole Method for the Laplace equation in three dimensions. Acta Numerica, 6:229– 269, 1997. CODEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Griewank:2003:MVA

[Gri03] Andreas Griewank. A mathematical view of automatic differentiation. Acta Numerica, 12:321–398, 2003. CODEN ANUMFU. ISBN 0-521-82523-

7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Giles:2002:AMP

[GS02] Michael B. Giles and Endre Süli. Adjoint methods for PDEs: a posteriori error analysis and postprocessing by duality. Acta Numerica, 11:145–236, 2002. CODEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (ebook). ISSN 0962-4929 (print), 1474-0508 (electronic).

Guttel:2017:NEP

[GT17] Stefan Güttel and Françoise Tisseur. The nonlinear eigenvalue problem. Acta Numerica, 26:1– 94, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
product/DE575DBE4F2A0A1D43E5E5F980FF65B3.

Gutknecht:1997:LTS

[Gut97] Martin H. Gutknecht. Lanczostype solvers for nonsymmetric linear systems of equations. Acta Numerica, 6:271–397, 1997. CO-DEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ghattas:2021:LPB

[GW21] Omar Ghattas and Karen Willcox. Learning physics-based models from data: perspectives from inverse problems and model reduction. Acta Numerica, 30:445–554, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
learning-physicsbased-models-
from-data-perspectives-from-
inverse-problems-and-model-
reduction/C072A4B417F8C3873ED75C1D63BBB31D.

Gunzburger:2014:SFE

[GWZ14] Max D. Gunzburger, Clayton G. Webster, and Guannan Zhang. Stochastic finite element methods for partial differential equations with random input data. Acta Numerica, 23:521– 650, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Gander:2025:TPH

[GWZ25] Martin J. Gander, Shu-Lin Wu, and Tao Zhou. Time parallelization for hyperbolic and parabolic problems. Acta Numerica, 34:385–489, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
time-parallelization-for-
hyperbolic-and-parabolic-
problems/E580B25A44766E6729A65D7AF05B1198.

Gander:2022:SMD

[GZ22a] Martin J. Gander and Hui Zhang. Schwarz methods by domain truncation. Acta Numerica, 31:1–134, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
schwarz-methods-by-domain-
truncation/4D988ACBCF39AC79F778227BE7028D04.

Geshkovski:2022:TOC

[GZ22b] Borjan Geshkovski and Enrique Zuazua. Turnpike in optimal control of PDEs, ResNets, and beyond. Acta Numerica, 31:135–263, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
turnpike-in-optimal-control-
of-pdes-resnets-and-beyond/
A5A8337380F0AC79BD20773659C5DFC8.

Hackbusch:2014:NTC

[Hac14] Wolfgang Hackbusch. Numerical tensor calculus. Acta Numerica, 23:651–742, 2014. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hagstrom:1999:RBC

[Hag99]  Thomas Hagstrom.  Radia-

tion boundary conditions for the numerical simulation of waves. Acta Numerica, 8:47–106, 1999. CODEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Higham:2010:CMF

[HAM10] Nicholas J. Higham and Awad H. Al-Mohy. Computing matrix functions. Acta Numerica, 19:159–208, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hegland:2001:DMT

[Heg01] Markus Hegland. Data mining techniques. Acta Numerica, 10:313–355, 2001. CODEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Henshaw:1996:AGG

[Hen96] William D. Henshaw. Automatic grid generation. Acta Numerica, 5:121–148, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Higham:1996:RAN

[Hig96] Nicholas J. Higham. Review of “Acta Numerica 1994 and Acta Numerica 1995, Cambridge University Press”. Mathematics Today, 32(1/2):28–??, 1996. ISSN 1361-2042.

Higdon:2006:NMO

[Hig06] Robert L. Higdon. Numerical modelling of ocean circulation. Acta Numerica, 15:385– 470, 2006. CODEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hiptmair:2002:FEC

[Hip02] R. Hiptmair. Finite elements in computational electromagnetism. Acta Numerica, 11:237– 339, 2002. CODEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Huan:2024:OED

[HJM24] Xun Huan, Jayanth Jagalur, and Youssef Marzouk. Optimal experimental design: Formulations and computations. Acta Numerica, 33:715–840, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
optimal-experimental-design-
formulations-and-computations/
38BBD0DC1A0386FDF306B6C0167DF7D9.

Heinkenschloss:2025:OPG

[HK25] Matthias Heinkenschloss and Drew P. Kouri. Optimization problems governed by systems of PDEs with uncertainties. Acta Numerica, 34:491–577, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
optimization-problems-governed-
by-systems-of-pdes-with-uncertainties/
AA58A4CF702C05F375F12242C8833C46.

Hairer:2003:GNI

[HLW03] Ernst Hairer, Christian Lubich, and Gerhard Wanner. Geometric numerical integration illustrated by the Störmer–Verlet method. Acta Numerica, 12:399– 450, 2003. CODEN ANUMFU. ISBN 0-521-82523-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Higham:2022:MPA

[HM22] Nicholas J. Higham and Theo Mary. Mixed precision algorithms in numerical linear algebra. Acta Numerica, 31:347–414, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        mixed-precision-algorithms-
                                        in-numerical-linear-algebra/
                                        43CA701BA29251B5790C653E66F46197.

Hochbruck:2010:EI

[HO10] Marlis Hochbruck and Alexander Ostermann. Exponential integrators. Acta Numerica, 19:209– 286, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hou:1995:NSF

[Hou95] Thomas Y. Hou. Numerical solutions to free boundary problems. Acta Numerica, 4:335–415, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hou:2009:BNB

[Hou09]  Thomas Y. Hou.  Blow-up

or no blow-up? A unified computational and analytic approach to 3D incompressible Euler and Navier–Stokes equations. Acta Numerica, 18:277– 346, 2009. CODEN ANUMFU. ISBN 0-521-19211-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Hesthaven:2022:RBM

[HPR22] Jan S. Hesthaven, Cecilia Pagliantini, and Gianluigi Rozza. Reduced basis methods for timedependent problems. Acta Numerica, 31:265–345, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
reduced-basis-methods-for-
timedependent-problems/B3A3572002FB0071F9E903AAEAB83087.

Han:2019:NAH

[HS19] Weimin Han and Mircea Sofonea. Numerical analysis of hemivariational inequalities in contact mechanics. Acta Numerica, 28:175–286, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
numerical-analysis-of-hemivariational-
inequalities-in-contact-mechanics/
FEE4F4CDD4F764901A99CCF6AF3A4D84.

Iserles:2000:LGM

[IMKNZ00] Arieh Iserles, Hans Z. Munthe-Kaas, Syvert P. Nørsett, and Antonella Zanna. Lie-group methods. Acta Numerica, 9:215–365, 2000. CODEN ANUMFU. ISBN 0-521-78037-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Ipsen:1998:RPR

[Ips98] Ilse C. F. Ipsen. Relative perturbation results for matrix eigenvalues and singular values. Acta Numerica, 7:151–201, 1998. CO-DEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Jin:2022:APS

[Jin22] Shi Jin. Asymptotic-preserving schemes for multiscale physical problems. Acta Numerica, 31:415–489, May 2022. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        asymptoticpreserving-schemes-
                                        for-multiscale-physical-problems/
                                        302602CD4C34A0A8E2ED6A3703D8D461.

Jin:2011:MCM

[JMS11] Shi Jin, Peter Markowich, and Christof Sparber. Mathematical and computational methods for semiclassical Schrödinger equations. Acta Numerica, 20:121– 209, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Kelley:2018:NMN

[Kel18] C. T. Kelley. Numerical methods for nonlinear equations. Acta Numerica, 27:207–287, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/

numerical-methods-for-nonlinear-
equations/B9E91D8ABB02ECC16B8715CAD263EA06.

Kreiss:1998:STD

[KL98] Heinz-Otto Kreiss and Jens Lorenz. Stability for timedependent differential equations. Acta Numerica, 7:203–285, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Kreiss:1992:PDT

[Kre92] Heinz-Otto Kreiss. Problems with different time scales. Acta Numerica, 1:101–139, 1992. CO-DEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Kuhn:2025:DRO

[KSW25] Daniel Kuhn, Soroosh Shafiee, and Wolfram Wiesemann. Distributionally robust optimization. Acta Numerica, 34: 579–804, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

URL https://www.cambridge.
org/core/journals/acta-numerica/
article/distributionally-
robust-optimization/5B4E65E3A5A2AEF24E218A6B34E6EAA2.

Kurganov:2018:FVS

[Kur18] Alexander Kurganov. Finitevolume schemes for shallowwater equations. Acta Numerica, 27:289–351, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

URL https://www.cambridge.
org/core/journals/acta-numerica/

                                        article/finitevolume-schemes-
                                        for-shallowwater-equations/
                                        AE2AC80D1E6E9F6BC0E68496A1C3EC52.

Kannan:2017:RAN

[KV17] Ravindran Kannan and Santosh Vempala. Randomized algorithms in numerical linear algebra. Acta Numerica, 26:95– 135, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/41CF2151FADE7757AA95C7FC15E43630.

Lasserre:2024:MSH

[Las24] Jean B. Lasserre. The moment-SOS hierarchy: Applications and related topics. Acta Numerica, 33:841–908, July 2024. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        momentsos-hierarchy-applications-
                                        and-related-topics/83B5483C595660346E598010DDC69200.

Lebrun:2012:SID

[LCBM12] M. Lebrun, M. Colom, A. Buades, and J. M. Morel. Secrets of image denoising cuisine. Acta Numerica, 21:475–576, 2012. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

LeBris:2005:CCP

[Le 05] Claude Le Bris. Computational chemistry from the perspective of numerical analysis. Acta Numerica, 14:363–444, 2005. CO-DEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

LeVeque:2011:TMA

[LGB11] Randall J. LeVeque, David L.

George, and Marsha J. Berger. Tsunami modelling with adaptively refined finite volume methods. Acta Numerica, 20:211–289, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Li:1997:NSM

[Li97] T. Y. Li. Numerical solution of multivariate polynomial systems by homotopy continuation methods. Acta Numerica, 6:399–436, 1997. CODEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Lim:2021:TC

[Lim21]  Lek-Heng Lim.  Tensors in

computations. Acta Numerica, 30:555–764, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
tensors-in-computations/F7EF66198DB10D3DE70B181D2E77D949.

Lasser:2020:CQD

[LL20] Caroline Lasser and Christian Lubich. Computing quantum dynamics in the semiclassical regime. Acta Numerica, 29:229–401, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/

journals/acta-numerica/article/
computing-quantum-dynamics-
in-the-semiclassical-regime/
05B21469F9D9557EA7A0117B6C7954E3.

Lin:2019:NMK

[LLY19] Lin Lin, Jianfeng Lu, and Lexing Ying. Numerical methods for Kohn–Sham density functional theory. Acta Numerica, 28:405–539, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
numerical-methods-for-kohnsham-
density-functional-theory/
755DFB88349DD5F1EE1E360AD61661BF.

LeFloch:2014:NMC

[LM14] Philippe G. LeFloch and Siddhartha Mishra. Numerical methods with controlled dissipation for small-scale dependent shocks. Acta Numerica, 23:743– 816, 2014. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Larson:2019:DFO

[LMW19] Jeffrey Larson, Matt Menickelly, and Stefan M. Wild. Derivativefree optimization methods. Acta Numerica, 28:287–404, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
derivativefree-optimization-
methods/84479E2B03A9BFFE0F9CD46CF9FCD289.

Lewis:1996:EO

[LO96] Adrian S. Lewis and Michael L.

Overton. Eigenvalue optimization. Acta Numerica, 5:149–190, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Luskin:2013:ACC

[LO13] Mitchell Luskin and Christoph Ortner. Atomistic-to-continuum coupling. Acta Numerica, 22:397–508, 2013. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Lelievre:2016:PDE

[LS16] Tony Lelièvre and Gabriel Stoltz. Partial differential equations and stochastic methods in molecular dynamics. Acta Numerica, 25:681–880, 2016. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Luskin:1996:CCM

[Lus96] Mitchell Luskin. On the computation of crystalline microstructure. Acta Numerica, 5:191–257, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Marz:1992:NMD

[Mär92]  Roswitha März.  Numerical

methods for differential algebraic equations. Acta Numerica, 1:141–198, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Mischaikow:2002:TTE

[Mis02] Konstantin Mischaikow. Topological techniques for efficient rigorous computation in dynamics. Acta Numerica, 11:435–477, 2002. CODEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (ebook). ISSN 0962-4929 (print), 1474-0508 (electronic).

Miekkala:1996:ISS

[MN96] Ulla Miekkala and Olavi Nevanlinna. Iterative solution of systems of linear differential equations. Acta Numerica, 5:259– 307, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

McLachlan:2002:SM

[MQ02] Robert I. McLachlan and G. Reinout W. Quispel. Splitting methods. Acta Numerica, 11:341–434, 2002. CO-DEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Meurant:2006:LCG

[MS06] Gérard Meurant and Zdeněk Strakoš. The Lanczos and conjugate gradient algorithms in finite precision arithmetic. Acta Numerica, 15:471–542, 2006. CO-DEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Morton:2007:FVM

[MS07] K. W. Morton and T. Sonar. Finite volume methods for hyperbolic conservation laws. Acta Numerica, 16:155–238, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Martinsson:2020:RNL

[MT20] Per-Gunnar Martinsson and Joel A. Tropp. Randomized numerical linear algebra: Foundations and algorithms. Acta Numerica, 29:403–572, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
randomized-numerical-linear-
algebra-foundations-and-algorithms/
4486926746CFF4547F42A2996C7DC09C.

Mehrmann:2023:CPH

[MU23] Volker Mehrmann and Benjamin Unger. Control of port-Hamiltonian differentialalgebraic systems and applications. Acta Numerica, 32:395–515, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
control-of-porthamiltonian-
differentialalgebraic-systems-
and-applications/102B84AD17404344FAC406DC2A40D284.

Marsden:2001:DMV

[MW01] J. E. Marsden and M. West.

Discrete mechanics and variational integrators. Acta Numerica, 10:357–514, 2001. CODEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Natterer:1999:NMT

[Nat99]  Frank Natterer.  Numerical

methods in tomography. Acta Numerica, 8:107–141, 1999. CO-DEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Neumaier:2004:CSC

[Neu04] Arnold Neumaier. Complete search in continuous global optimization and constraint satisfaction. Acta Numerica, 13:271– 369, 2004. CODEN ANUMFU. ISBN 0-521-83811-8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Nesterov:2013:NNM

[NN13] Yurii Nesterov and Arkadi Nemirovski. On first-order algorithms for l_{1} /nuclear norm minimization. Acta Numerica, 22:509–575, 2013. CO-DEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Nocedal:1992:TAU

[Noc92] Jorge Nocedal. Theory of algorithms for unconstrained optimization. Acta Numerica, 1:199–242, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Neunzert:1995:PMB

[NS95] Helmut Neunzert and Jens Struckmeier. Particle methods for the Boltzmann equation. Acta Numerica, 4:417–457, 1995. CODEN ANUMFU. ISBN 0-521-48255-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Neilan:2017:NAS

[NSZ17] Michael Neilan, Abner J. Salgado, and Wujun Zhang. Numerical analysis of strongly nonlinear PDEs. Acta Numerica, 26:137–303, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

URL https://www.cambridge.
org/core/product/CB53724C153D209910B6EAA818E6B76D.

Nemirovski:2008:IPM

[NT08] Arkadi S. Nemirovski and Michael J. Todd. Interior-point methods for optimization. Acta Numerica, 17:191–234, 2008. CODEN ANUMFU. ISBN 0-521-51642-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Oden:2018:AMP

[Ode18] J. Tinsley Oden. Adaptive multiscale predictive modelling. Acta Numerica, 27:353–450, 2018. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
adaptive-multiscale-predictive-
modelling/23390D2DBEFCB051ABA6A1A854260DB9.

Olver:2020:FAU

[OST20] Sheehan Olver, Richard Mikaël Slevinsky, and Alex Townsend. Fast algorithms using orthogonal polynomials. Acta Numerica, 29:573–699, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        fast-algorithms-using-orthogonal-
                                        polynomials/4FAD8C7C28EC20EE7583465C1A89AA3D.

Ozyesil:2017:SSM

[ÖVBS17] Onur Özyesil, Vladislav Voroninski, Ronen Basri, and Amit Singer. A survey of structure from motion. Acta Numerica, 26:305–364, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/C4B2E7BB10BC2C11AF71BC80B584D378.

Parlett:1995:NA

[Par95] Beresford N. Parlett. The new qd algorithms. Acta Numerica, 4:459–491, 1995. CODEN ANUMFU. ISBN 0-521-48255-

0. ISSN 0962-4929 (print), 1474-

0508 (electronic).

Peskin:2002:IBM

[Pes02] Charles S. Peskin. The immersed boundary method. Acta Numerica, 11:479–517, 2002. CO-DEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Pothen:2019:AAC

[PFM19] Alex Pothen, S. M. Ferdous, and Fredrik Manne. Approximation algorithms in combinatorial scientific computing. Acta Numerica, 28:541–633, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
approximation-algorithms-
in-combinatorial-scientific-
computing/8B1B0396F6DF012CFBDF52E686197DFC.

Pinkus:1999:ATM

[Pin99] Allan Pinkus. Approximation theory of the MLP model in neural networks. Acta Numerica, 8:143–195, 1999. CODEN ANUMFU. ISBN 0-521-77088-

2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Pironneau:2012:PFD

[Pir12] Olivier Pironneau. Pricing futures by deterministic methods. Acta Numerica, 21:577– 671, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Petzold:1997:NSH

[PJY97] Linda R. Petzold, Laurent O.

Jay, and Jeng Yen. Numerical solution of highly oscillatory ordinary differential equations. Acta Numerica, 6:437–483, 1997. CO-DEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Platen:1999:INM

[Pla99] Eckhard Platen. An introduction to numerical methods for stochastic differential equations. Acta Numerica, 8:197–246, 1999.

CODEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Powell:1998:DSA

[Pow98] M. J. D. Powell. Direct search algorithms for optimization calculations. Acta Numerica, 7:287– 336, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Quarteroni:2017:CSM

[QMV17] A. Quarteroni, A. Manzoni, and C. Vergara. The cardiovascular system: Mathematical modelling, numerical algorithms and clinical applications. Acta Numerica, 26:365– 590, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        product/B79D5D7B17499F8758150FEEC4207916.

Reich:2019:DAS

[Rei19]  Sebastian Reich.  Data as-

similation: The Schrödinger perspective. Acta Numerica, 28:635–711, May 01, 2019. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        data-assimilation-the-schrodinger-
                                        perspective/9C27410C665B9676110813C31CD53446.

Ringhofer:1997:CMS

[Rin97] Christian Ringhofer. Computational methods for semiclassical and quantum transport in semiconductor devices. Acta Numerica, 6:485–521, 1997. CODEN ANUMFU. ISBN 0-521-59106-

6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Rump:2010:VMR

[Rum10] Siegfried M. Rump. Verification methods: Rigorous results using floating-point arithmetic. Acta Numerica, 19:287– 449, 2010. CODEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (elec-

tronic). URL http://www.ti3.
tu-harburg.de/paper/rump/
Ru10.pdf.

Saad:2025:AMF

[Saa25] Yousef Saad. Acceleration methods for fixed-point iterations. Acta Numerica, 34:805–890, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
acceleration-methods-for-
fixedpoint-iterations/2FABE7BD50568960C8CD86A6F4D77D72.

Sabin:1994:NGS

[Sab94] Malcolm Sabin. Numerical geometry of surfaces. Acta Numerica, 3:411–466, 1994. CODEN ANUMFU. ISBN 0-521-46181-

2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Sethian:1996:TAA

[Set96] James A. Sethian. Theory, algorithms, and applications of level set methods for propagating interfaces. Acta Numerica, 5:309– 395, 1996. CODEN ANUMFU. ISBN 0-521-57234-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Schwab:2011:STD

[SG11] Christoph Schwab and Claude Jeffrey Gittelson. Sparse tensor discretizations of high-dimensional parametric and stochastic PDEs. Acta Numerica, 20:291–467, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Shapiro:2007:SAG

[Sha07] Vadim Shapiro. Semi-analytic geometry with R-functions. Acta Numerica, 16:239–303, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Shu:2020:ENO

[Shu20] Chi-Wang Shu. Essentially nonoscillatory and weighted essentially non-oscillatory schemes. Acta Numerica, 29:701–762, May 2020. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
essentially-nonoscillatory-
and-weighted-essentially-
nonoscillatory-schemes/28F4F341659059051F12A4CC7924905C.

Schutte:2023:OTB

[SKH23] Christof Schütte, Stefan Klus, and Carsten Hartmann. Overcoming the timescale barrier in molecular dynamics: Transfer operators, variational principles and machine learning. Acta Numerica, 32:517–673, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
overcoming-the-timescale-
barrier-in-molecular-dynamics-
transfer-operators-variational-
principles-and-machine-learning/
AFE1143DC57B9D2B33774997A620E277.

Sloan:1992:EAB

[Slo92] Ian H. Sloan. Error analysis of boundary integral methods. Acta Numerica, 1:287–339, 1992. CO-DEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Smale:1997:CTN

[Sma97] Steve Smale. Complexity theory and numerical analysis. Acta Numerica, 6:523–551, 1997. CO-DEN ANUMFU. ISBN 0-521-59106-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Smith:2004:MCM

[SNCH04] N. P. Smith, D. P. Nickerson, E. J. Crampin, and P. J. Hunter. Multiscale computational modelling of the heart. Acta Numerica, 13:371–431, 2004. CODEN ANUMFU. ISBN 0-521-83811-

8. ISSN 0962-4929 (print), 1474-0508 (electronic).

Sorensen:2002:NML

[Sor02] Danny C. Sorensen. Numerical methods for large eigenvalue problems. Acta Numerica, 11:519–584, 2002. CO-DEN ANUMFU. ISBN 0-521-81876-1, 0-511-55014-6 (e-book). ISSN 0962-4929 (print), 1474-0508 (electronic).

Sanz-Serna:1992:SIH

[SS92] J. M. Sanz-Serna. Symplectic integrators for Hamiltonian problems: an overview. Acta Numerica, 1:243–286, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

Scott:2025:SLL

[ST25] Jennifer Scott and Miroslav Tuma. Sparse linear leastsquares problems. Acta Numerica, 34:891–1010, July 2025. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

                                        //www.cambridge.org/core/
                                        journals/acta-numerica/article/
                                        sparse-linear-leastsquares-
                                        problems/F992BEAE18E16437E680624619C79DE6.

Stuart:1994:NAD

[Stu94] Andrew M. Stuart. Numerical analysis of dynamical systems. Acta Numerica, 3:467–572, 1994. CODEN ANUMFU. ISBN 0-521-46181-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Stuart:2010:IPB

[Stu10] A. M. Stuart. Inverse problems: a Bayesian perspective. Acta Numerica, 19:451–559, 2010. CO-DEN ANUMFU. ISBN 0-521-19284-6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Stynes:2005:SSC

[Sty05] Martin Stynes. Steady-state convection-diffusion problems. Acta Numerica, 14:445–508, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Schaback:2006:KTM

[SW06] Robert Schaback and Holger Wendland. Kernel techniques: from machine learning to meshless methods. Acta Numerica, 15:543–639, 2006. CODEN ANUMFU. ISBN 0-521-86815-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Tadmor:2003:EST

[Tad03] Eitan Tadmor. Entropy stability theory for difference approximations of nonlinear conservation laws and related timedependent problems. Acta Numerica, 12:451–512, 2003. CO-DEN ANUMFU. ISBN 0-521-82523-7. ISSN 0962-4929 (print), 1474-0508 (electronic).

Tadmor:2007:FMC

[Tad07] Eitan Tadmor. Filters, mollifiers and the computation of the Gibbs phenomenon. Acta Numerica, 16:305–378, 2007. CO-DEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Temme:2007:NAS

[Tem07]  Nico M. Temme.  Numer-

ical aspects of special functions. Acta Numerica, 16:379– 478, 2007. CODEN ANUMFU. ISBN 0-521-87743-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Temlyakov:2008:GA

[Tem08] V. N. Temlyakov. Greedy approximation. Acta Numerica, 17:235–409, 2008. CODEN ANUMFU. ISBN 0-521-51642-

0. ISSN 0962-4929 (print), 1474-

0508 (electronic).

Tsai:2005:TVL

[TO05] Yen-Hsi Richard Tsai and Stanley Osher. Total variation and level set methods in image science. Acta Numerica, 14:509– 573, 2005. CODEN ANUMFU. ISBN 0-521-85807-0. ISSN 0962-4929 (print), 1474-0508 (electronic).

Todd:2001:SO

[Tod01]  M. J. Todd.  Semidefinite

optimization. Acta Numerica, 10:515–560, 2001. CO-DEN ANUMFU. ISBN 0-521-80312-8. ISSN 0962-4929 (print), 1474-0508 (elec-

                                        tronic).  URL http://www.
                                        orie.cornell.edu/~miketodd/
                                        soa4.ps.Z;   http://www.orie.
                                        cornell.edu/~miketodd/todd.
                                        html;   http://www.orie.cornell.
                                        edu/~miketodd/soa5.ps.

Trefethen:1999:CP

[Tre99] Lloyd N. Trefethen. Computation of pseudospectra. Acta Numerica, 8:247–295, 1999. CO-DEN ANUMFU. ISBN 0-521-77088-2. ISSN 0962-4929 (print), 1474-0508 (electronic).

Tuncel:2023:LOH

[TV23] Levent Tunçel and Lieven Vandenberghe. Linear optimization over homogeneous matrix cones. Acta Numerica, 32:675–747, May 2023. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
linear-optimization-over-
homogeneous-matrix-cones/
0033A4A37E35CFF3BC89803365848196.

vanDorsselaer:1993:LSA

[vDKS93] J. L. M. van Dorsselaer, J. F.

B. M. Kraaijevanger, and M. N. Spijker. Linear stability analysis in the numerical solution of initial value problems. Acta Numerica, 2:199–237, 1993. CO-DEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Wanner:2010:KNN

[Wan10] G. Wanner. Kepler, Newton and numerical analysis. Acta Numerica, 19:561–598, 2010. CODEN ANUMFU. ISBN 0-521-19284-

6. ISSN 0962-4929 (print), 1474-0508 (electronic).

Watson:1998:CND

[Wat98] G. A. Watson. Choice of norms for data fitting and function approximation. Acta Numerica, 7:337–377, 1998. CODEN ANUMFU. ISBN 0-521-64316-3. ISSN 0962-4929 (print), 1474-0508 (electronic).

Wathen:2015:P

[Wat15] A. J. Wathen. Preconditioning. Acta Numerica, 24:329– 376, 2015. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Wohlmuth:2011:VCD

[Woh11]  Barbara Wohlmuth.  Varia-

tionally consistent discretization schemes and numerical algorithms for contact problems. Acta Numerica, 20:569– 734, 2011. CODEN ANUMFU. ISBN 1-107-01086-1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Wright:1992:IMC

[Wri92]  Margaret H. Wright.  Inte-

rior methods for constrained optimization. Acta Numerica, 1:341–407, 1992. CODEN ANUMFU. ISBN 0-521-41026-6, 0-521-42291-4. ISSN 0962-4929 (print), 1474-0508 (electronic).

                                        URL ftp://netlib.att.com/
                                        netlib/att/cs/doc/91/4-10.
                                        ps.Z.

Wampler:2011:NAG

[WS11] Charles W. Wampler and Andrew J. Sommese. Numerical algebraic geometry and algebraic kinematics. Acta Numerica, 20:469–567, 2011. CODEN ANUMFU. ISBN 1-107-01086-

1. ISSN 0962-4929 (print), 1474-0508 (electronic).

Wang:2021:MCL

[WZZ21] Wei Wang, Lei Zhang, and Pingwen Zhang. Modelling and computation of liquid crystals. Acta Numerica, 30:765–851, May 2021. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
journals/acta-numerica/article/
modelling-and-computation-
of-liquid-crystals/DFF15BBCE2D4B5773625923445CEAE45.

Xu:2017:AMM

[XZ17] Jinchao Xu and Ludmil Zikatanov.

Algebraic multigrid methods. Acta Numerica, 26:591–721, 2017. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic). URL https:

//www.cambridge.org/core/
product/8FFBCDA39DB9631667396C9CD1F223BF.

Yserentant:1993:ONC

[Yse93] Harry Yserentant. Old and new convergence proofs for multigrid methods. Acta Numerica, 2:285– 326, 1993. CODEN ANUMFU. ISBN 0-521-44356-3. ISSN 0962-4929 (print), 1474-0508 (electronic).